The number of books N to be sold at a bookfare is given by the function
N(p) = 25000 / (3p + k)
where p is the price per book and k is a constant. If 1000 books are sold at $7 per book, how many books will be sold at the price of $10 per book?

Answer by nerdybill(7384)

The number of books N to be sold at a bookfare is given by the function
N(p) = 25000 / (3p + k)
where p is the price per book and k is a constant. If 1000 books are sold at $7 per book, how many books will be sold at the price of $10 per book?
.
First, use: "1000 books are sold at $7 per book" to find k:
N(p) = 25000 / (3p + k)
1000 = 25000 / (3(7) + k)
1000 = 25000 / (21 + k)
1000(21 + k) = 25000
21000 + 1000k = 25000
1000k = 4000
k = 4
.
Our formula is now:
N(p) = 25000 / (3p + 4)
Now we can answer: "how many books will be sold at the price of $10 per book?"
N(p) = 25000 / (3(10) + 4)
N(p) = 25000 / (30 + 4)
N(p) = 25000/34
N(p) = 735.29
But, since we can't sell a partial book:
N(p) = 735